jhonyy9 (jhonyy9):

let formula 2^n *(2m +1) with n and m natural numbers plus/(with)* zero * - because zero not accepted natural - using this formula can be writing ALL numbers ,like for example : 2^0 *(2*0 +1) = 1 2^1 *(2*0 +1) = 2 2^0 *(2*1 +1) = 3 2^2 *(2*0 +1) = 4 ..... can being this proven so how ?

ganeshie8 (ganeshie8):

`n=0` gives all odd natural numbers since every odd number can be expressed as `2m+1` by division algorithm

jhonyy9 (jhonyy9):

yes thank you so but in case of even numbers ?

jhonyy9 (jhonyy9):

2^1 *(2*0 +1)= 2 2^2 *(2*0 +1)= 4 2^3 *(2*0 +1)= 8 .... - so we get every power of 2

myininaya (myininaya):

what are you trying to prove?

jhonyy9 (jhonyy9):

that using this formula we can getting every natural numbers

OpenStudy (perl):

you are trying to prove can any natural number be written as the product of a power of 2 and an odd number

OpenStudy (perl):

You want to prove that for any integer k, there exist m, n such that 2^n *(2m +1)= k

OpenStudy (perl):

2 = 2^1 * (2*0 + 1)

OpenStudy (perl):

every odd integer can be written as 2m+1

OpenStudy (perl):

so we only have to look at even numbers

OpenStudy (perl):

show that for any even number a=2t , a = 2^n ( 2m + 1)

OpenStudy (perl):

if a = 2t, then t is either even or odd

OpenStudy (perl):

case 1: t is even, then t = 2h for some integer h 2 (2h) = 2^n (2m+1)

OpenStudy (perl):

at some point h as to be odd

OpenStudy (ikram002p):

2^n *(2m +1) from D.A 2m+1 is the formula of all odd numbers , thus 2^0 *(2m +1)=2m+1 gives all odd u cant represent odd numbers if \(n\neq 0\) now let n be any number n>0 2^n *(2m +1) = 2 (2^n-1 *(2m +1) = 2k all even numbers :)

OpenStudy (ikram002p):

here 1 case to get even numbers :- let n=1 , m>=0 2^n *(2m +1) 2(1)=2 2(3)=6 2(4)=8 2(5)=10 ..... and let n=2 , m=0 to get 4 4(1)=0 ...

ganeshie8 (ganeshie8):

clever !

jhonyy9 (jhonyy9):

@ikram002p you have wrote 2(1)=2 this mean m=0 and n=1 so on next step 2(3)=6 so in this case n bot equal 3 and m=0 ? on the second case 2(3)= 6 not will be the right result 8 because 2^3 =8 ? how you think it please ?

ganeshie8 (ganeshie8):

thats a good catch

OpenStudy (ikram002p):

2(3) n=0,m=1 :O 2(5) n=0 , m=2 ... and go on with n=0 , m>=0 condition

jhonyy9 (jhonyy9):

ok but you have wrote in the first case 2(1)=2

OpenStudy (ikram002p):

im i missing something in the question ? i can chose m and n as i want , right ?

ganeshie8 (ganeshie8):

yes but how do u construct the even number 8 using n=1 and m>= 0 ?

OpenStudy (ikram002p):

oh sorry my bad xD i made a typo for 8

OpenStudy (ikram002p):

ok then 3 cases for numbers :- 1- n =0 , and m>=0 we get all odds 2- m=0 , and n>0 we get numbers n s.t n=2^k 3- n=1 , and m>=0 we get all even number n s.t n= 2k , and k has an odd factor

ganeshie8 (ganeshie8):

BINGO!

OpenStudy (ikram002p):

:3

jhonyy9 (jhonyy9):

ty @ganeshie8 and @ikram002p ! so van being this proven by mathinduction or any ideas ?

jhonyy9 (jhonyy9):

can being sorry

ganeshie8 (ganeshie8):

sure, ikram u want to try proving it by induction?

OpenStudy (ikram002p):

hmm we used proof by division algorithm :) we give cases

OpenStudy (ikram002p):

sure its trivial :)

ganeshie8 (ganeshie8):

i feel proving it by contradiction is more natural here though

OpenStudy (ikram002p):

as a start let k=2^n *(2m +1) then show that k+1 can be written as 2^j*(2i+1)

OpenStudy (ikram002p):

all works

jhonyy9 (jhonyy9):

@ganeshie8 above you have wrote ba contradiction mean by reductio ad absurdum ?

jhonyy9 (jhonyy9):

by contradiction

ganeshie8 (ganeshie8):

Exactly! we have already established that odd numbers can be constructed by division alogirthm. For even numbers, we can use contradiction

ganeshie8 (ganeshie8):

suppose there exists some even number which cannot be represented in form : \(2^n(2m+1)\)

ganeshie8 (ganeshie8):

by definition an even number is divisible by 2. that means an even number can be divided repeateadly by 2 till you get an odd number.

ganeshie8 (ganeshie8):

you stop dividing once you reach an odd number. after repeated divisions the original even number number can be represented as \(2^{n}(2m+1)\)

ganeshie8 (ganeshie8):

actually this argument looks nice for a direct proof :)

ganeshie8 (ganeshie8):

do u have a proof using induction @jhonyy9 ?

ganeshie8 (ganeshie8):

it would be interesing to see how induction works here xD

OpenStudy (ikram002p):

even for induction , we would also have to discuss 3 cases of the number ;)

OpenStudy (ikram002p):

when k is odd k=2m+1 for some m in N k+1 =2m+2=2(m+1) m+1 might be even or odd, ... continue like this so u would conclude the other two cases

jhonyy9 (jhonyy9):

one idea for even numbers 2^n *(2m +1) = 2(2^(n-1) *(2m +1) opinions please ?

jhonyy9 (jhonyy9):

so for n=N and m=0 will result all power of 2

OpenStudy (anonymous):

so wat Nnesha :O